# DeepMIDE: A Multivariate Spatio-Temporal Method for Ultra-Scale Offshore Wind Energy Forecasting

### v1

@@ 5.1 Forecasting results
Figure 5(a) shows the spatially averaged mean absolute error (MAE) values for all models plotted against the forecast horizon (in hours) for the three heights considered, namely: $100$m, $140$m, and $180$m. Specifically, the MAE for a representative benchmark $\mathcal{B}$, denoted by $\text{MAE}^{(g)}_{h}(\mathcal{B})$, is calculated as in (18).
@@
where ${Z^{(g)}_{T+h,i}(\mathbf{s}_{j})}$ and $\hat{Z}^{(g)}_{T+h,i}(\mathbf{s}_{j})$ denote the actual wind speed and forecasts at location $\mathbf{s}_{j}$ and height $g$ for forecasting horizon $h$ and $i$th forecasting roll, respectively. To provide a more condensed view of the results, Figure 5(b) shows the correspondent boxplots of the percentage reduction in MAE across all locations. The percentage reduction is computed using the average of all absolute errors taken across all forecast rolls and locations for the two methods being compared, as expressed in (19).
$\text{IMP}(\mathcal{B}^{*},\mathcal{B})=100\times(1-\frac{\text{MAE}^{(g)}_{h}(\mathcal{B}^{*})}{\text{MAE}^{(g)}_{h}(\mathcal{B})}),$ (19)
where $\mathcal{B}^{*}$ is the proposed DeepMIDE model, while $\mathcal{B}$ is any other benchmark including PER, ARIMA, STGP, DeepAR. A first glance to Figure 5 shows that, except for few instances, DeepMIDE consistently outperforms all benchmarks across all horizons and heights. DeepMIDE is typically followed by DeepAR as a fairly distant second (average percentage reduction in MAE of $3.26$%). PER and ARIMA perform fairly similarly ($3.91$% and $4.10$% percentage reductions in MAE, respectively). Trailing behind is STGP with an average percent reduction in forecast error of $8.16$% in favor of DeepMIDE.
Figure 5: (a) Average MAE values (aggregated in hourly intervals) versus the forecast horizon for all models, at different heights, namely: $100$m, $140$m, and $180$m. (b) Boxplots of the percentage reduction in MAE, across all rolls and forecast horizon, grouped by height.
Surprisingly, we find that STGP demonstrated weaker performance relative to PER and ARIMA. Upon further investigation, we found that STGP performed fairly well in short-term horizons, but suffered the most in longer-term extrapolation (beyond $\sim$6666 hours ahead), primarily due to the tendency of GP-based models to revert to their mean functions as the spatial and temporal correlations gradually die out. Notably, DeepMIDE achieved the best forecast performance for the first hour (average MAE of $0.660$), compared to an MAE of $0.686$ for ARIMA, $0.687$ for STGP, and $0.689$ for PER. Importantly, we observe that the improvement from DeepMIDE compared to the benchmarks increases as the forecast horizon becomes longer.
We also find that, on average, all models (including DeepMIDE) have slightly higher MAE values at higher altitudes. However, this is likely an artifact of the wind speeds being stronger as the altitude increases. When compared to PER and ARIMA, the improvement of DeepMIDE appears to be fairly consistent across all heights (in the range of $\sim$4444-$5$reduction in MAE). For STGP, the improvement is maximal at the $180$-m height, whereas for DeepAR, the opposite is observed. Thus, we do not necessarily see a clear trend in how DeepMIDE performs across different heights compared to the benchmark models.
Figure 6(a) shows the hourly MAE values $\text{MAE}^{(\mathbf{s})}_{h}(\mathcal{B})$ (averaged over height) plotted against the forecast horizon, across the three spatial sites considered: E05N, E06, and ASOW6, which is calculated as in (20).
@@
Figure 6(b) shows the corresponding boxplots of the percentage MAE reduction. The results are largely similar to those presented in Figure 5. However, this time, there is a clear trend in how DeepMIDE performs across different spatial locations, relative to the baseline models. Specifically, the maximal improvement appears at E06 (in the range of $\sim$5555-$10$reduction in MAE). We postulate that this may be due to E06 being at a central location, in between the two other spatial sites, making the gain from including neighborhood information greater. E06 is followed by E05N ($\sim$3333-$8$reduction in MAE) where its position as a downstream location appears to benefit from the advection dynamics modeled in DeepMIDE. In contrast, ASOW6, being both the southernmost and an upstream location, exhibits the smallest improvement from DeepMIDE ($\sim$2222-$7$reduction in MAE).
Figure 6: Average MAE values (aggregated in hourly intervals) versus the forecast horizon for all models, at different locations, namely: E05N, E06, and ASOW6. (b) Boxplots of the percentage reduction in MAE, across all rolls and forecast horizon, grouped by spatial location.
@@
Figure 7 shows the time series of the one-step wind speed forecasts from DeepMIDE on top of the actual measurements at the hub height of $140$m, during a select time interval in June 2021. The time series of the forecasts show faithful agreement with the actual measurements for all three locations. The corresponding $95$% forecast intervals visually show satisfactory coverage and sharpness. This probabilistic nature of IDE-based models make them well-suited to inform subsequent operational decision-making under uncertainty for offshore wind farms (Pinson, 2013; Papadopoulos et al., 2024a).
@@
Figure 7: Time series of $1$-step ahead forecasts from DeepMIDE on top of the actual wind speed observations at the hub height of $140$m for all three locations ($95$% forecast intervals shown).
@@
Wind speed is the major determinant of wind power, and hence, we would like to illustrate the value of our approach for improving wind power prediction accuracy. To do so, we need to convert the wind speed forecasts into the wind power domain. As there are no operational wind farms in the NY/NJ Bight region yet, we leverage an external set of operational data from a wind farm in the United States to construct a statistical wind power curve which will be used for wind-to-power conversion (Ding, 2019). The power curve is multivariate and it takes two inputs: the hub-height wind speed and the vertical wind shear. The latter reflects the vertical change in wind speed. Let the wind shear be denoted by $S$, then an estimate for wind shear is computed as in (21).
$S=\frac{\log{(Z_{2}/Z_{1})}}{\log{(h_{1}/h_{2})}},$ (21)
@@
where $Z_{2}$ and $Z_{1}$ are the wind speeds measured at heights $h_{1}$ and $h_{2}$, respectively (Lee et al., 2015). For our data, wind speeds are measured at heights of $100$m, $140$m, and $180$m, with $140$m considered as the hub-height. Consequently, two wind shear values are calculated: an above-hub wind shear ($S_{a}$) using the $180$m/$140$m pair, and a below-hub wind shear ($S_{b}$) using the $140$m/$100$m pair. Figure 8(a) visualizes the fitted wind power curve (using a multivariate GP model akin to the one proposed by Golparvar et al. (2021)) with hub height wind speed and above-hub wind shear $S_{a}$. The power output is scaled to the $[0,100]$ interval, where $100$ represents the maximum rated capacity. By applying the constructed power curve, we transform the wind speed observations, along with their forecasts from the four competing models into corresponding wind power values.
@@
Figure 8(b) presents the boxplots of the improvement across all forecasting rolls, horizons, locations, and heights in wind power. DeepMIDE significantly outperforms all its competitors, demonstrating an average improvement of $8.89$% in wind power forecasting compared to other benchmarks. Enhanced accuracy in wind speed forecasts directly contributes to greater improvements in wind power predictions. Theoretically, wind speed is related to wind power through a cubic relationship, and hence, large improvements in forecasting wind speed are expected to yield even greater improvements in predicting wind power.
@@ 5.2 Discussion
To further elucidate the inner workings of DeepMIDE, we take a closer look at the advection parameter estimates, which are produced using the deep-learning-based functional mapping. Let the norm of the estimated advection vectors at height $g$ and time $t$ denote as $\Lambda_{t}^{g}=\|\boldsymbol{\theta}_{t}^{g}\|$. Figure 9 shows the time series plot of $\Lambda_{t}^{2}$ at height $140$m (corresponding to height index $g=2$ in our study), clearly illustrating a dynamic pattern over time. We zoom in on two consecutive NWP wind speed maps from 21-Jun-2021 02:00:00 to 03:00:00 (orange rectangle, left) and from 04-Jul-2021 05:00:00 to 06:00:00 (green rectangle, right). During the first time interval, the wind speed maps visually show strong advection effects over time, which appears to be captured by the model as reflected by the relatively large advection parameter predictions in the corresponding time series. In contrast, the second time interval shows significantly weaker advection effects (with the two consecutive images being almost static), with correspondingly smaller predicted values in the advection parameter time series. These two examples demonstrate how the deep learning model is able to translate the information from those maps into physically meaningful estimates of advection. Figure 10 shows the boxplots of all $\{\Lambda_{t}^{g}\}_{t=1}^{T}$, partitioned by wind speed (x-axis) and height (different colors), which clearly shows that the values of the predicted advection parameters increase at greater heights and stronger winds—a conclusion which we have observed in our preliminary data analysis of Section 3.3. This further confirms the ability of DeepMIDE to capture real-world aspects of the data.
Figure 9: Time series plot showing the norm of the predicted advection parameters at height $140$m, alongside spatial wind speed images from the numerical model at three consecutive time steps, for two different time intervals: 21-Jun-2021 02:00:00 to 03:00:00 (orange rectangle, left) and from 04-Jul-2021 05:00:00 to 06:00:00 (green rectangle, right).
@@ 6 Conclusions
As the upscaling of offshore wind turbines to unprecedented heights continues, multi-height wind speed forecasts across the now-much-larger rotor area of those ultra-scale generators are now needed. In this work, we propose the first attempt for multi-height spatio-temporal wind energy forecasting. Our model, dubbed DeepMIDE, is formulated as a multivariate integro-difference equation model, which fully embraces the full vertical wind profile by incorporating spatial, temporal, and height dependencies, in order to reflect the true complexity of wind interactions coincident with ultra-scale offshore wind turbines. An embedded deep learning architecture directly acts on streams of high-dimensional exogenous weather information in order to encode the physics of local wind field formation in a low-dimensional set of information-rich advection vectors. Those vectors, along with other parameters, are then plugged back into the statistical model for probabilistic space-time forecasting. The proposed approach marries the probabilistic rigor and parsimony of the multivariate integro-difference equation framework, with the representative power of the advanced deep transformer architecture which are used as a “physics extractor” within the statistical model. Evaluated using real-world data from the U.S. Mid Atlantic, our model demonstrates significant improvements in both wind speed and power forecasting compared with prevalent forecasting benchmarks.
Future research will investigate more sophisticated modeling structures for the latent process model within the MIDE framework. This exploration may include handling non-Gaussian data and non-linear dependence, which would obviate the need for data transformations or linear assumptions, but would likely complicate the parameter inference process.
@@
The supplementary materials appended to this manuscript include: (1) Details of the architecture and implementation of the deep learning model shown in Figure 4; (2) Access to codes and data used to reproduce the results of DeepMIDE.

### v2

@@ 5.1 Forecasting results
[…]
@@
where $Z^{(g)}_{T+h,i}(\mathbf{s}_{j})$} and $\hat{Z}^{(g)}_{T+h,i}(\mathbf{s}_{j})$ denote the actual wind speed and forecasts at location $\mathbf{s}_{j}$ and height $g$ for forecasting horizon $h$ and $i$th forecasting roll, respectively. The percentage reduction in forecast error is computed using the average of all absolute errors taken across all forecast rolls and locations for the two methods being compared, as expressed in (22).
[…]
where $\mathcal{B}^{*}$ is the proposed DeepMIDE model, while $\mathcal{B}$ is any benchmark model. Figure 5(a) shows the MAE across all locations and heights (We exclude NWP due to its poor predictive performance). A first glance at Table 1 and Figure 5(a) shows that DeepMIDE consistently outperforms all benchmarks across all horizons and heights (100m, 140m, and 180m), with improvements ranging from approximately 3.63.6% to 4.84.8% compared to other statistical models (PER, ARIMA, STGP), and about 3.93.9% to 7.27.2% compared to statistical deep learning models (DeepAR and DeepIDE-H). The physics-based NWP forecasts from RU-WRF exhibit significantly higher MAE. This is unsurprising considering that NWP models are not competitive at these short look-ahead horizons (ye2023ieee). Despite that, NWP data remains valuable in conveying physically meaningful information about advection dynamics, which are harnessed by DeepMIDE’s embedded deep learning architecture.
It is also worth noting that, at the $30$-minute horizon ($0.5$ hours ahead), where it is generally hard to beat simpler statistical models such as persistence and autoregressive methods, DeepMIDE maintains a small but meaningful margin of improvement, indicating it can capture immediate wind fluctuations, by virtue of its dynamical statistical representation. As the lead time increases, all models experience some degradation in accuracy, but DeepMIDE’s error grows more slowly than other benchmarks. DeepIDE-H seems to overtake PER as the second-best performing model as the forecast horizon extends ($h>1$ hour ahead). Yet, it is still outperformed by DeepMIDE as clearly shown in Figure 5(a), further confirming the merit of the multi-output approach pursued herein. At the $2$-hour ahead horizon, DeepMIDE still maintains a noticeably lower MAE than any other method, while the error of the DeepAR, for example, increases more rapidly with forecast horizon. DeepAR, although also a statistical deep learning model, shows lower performance. Unlike the proposed DeepMIDE, which is fundamentally a statistical model augmented by deep learning as a “physics extractor”, DeepAR is fundamentally a deep learning-based model and may require more training data to get better performance. Trailing behind is STGP, which does not appear to be capable of matching the performance of simpler statistical models in very short-term horizons (namely, PER and ARIMA), nor the more sophisticated Deep IDE models at longer horizons. Yet, STGP still outperforms the DeepAR and NWP models, likely due to its expressive kernel structure shown in (19).
Another interesting observation from Table 1 is that DeepMIDE achieves its greatest improvements at the 140m height compared to other benchmarks suggesting an advantage in leveraging central heights by effectively capturing advection and vertical cross-dependencies. We also find that, on average, all models (including DeepMIDE) have slightly higher MAE values at higher altitudes. However, this is likely an artifact of the wind speeds being stronger as the altitude increases.
Table 1: Average MAE values (aggregated in 30-min intervals) versus the forecast horizon $h$ for all models, at different heights, namely: 100100m, 140140m, and 180180m. Bold-faced and underlined values denote the best and second-best performing models, respectively.
Physics-Based Statistical Methods Statistical Deep Learning Methods
Location $h$ (hour) NWP PER ARIMA STGP DeepAR DeepIDE-H DeepMIDE
0.5 1.644 0.504 0.503 0.504 0.510 0.541 0.497
1 1.459 0.844 0.848 0.848 0.882 0.874 0.816
100mm 1.5 1.453 1.092 1.102 1.103 1.144 1.084 1.046
2 1.412 1.338 1.351 1.354 1.389 1.289 1.280
Avg 1.492 0.944 0.951 0.952 0.981 0.947 0.910
IMP($\%$) 39.008 3.602 4.311 4.412 7.238 3.907
0.5 1.751 0.507 0.511 0.507 0.515 0.543 0.499
1 1.556 0.860 0.865 0.866 0.889 0.890 0.826
140mm 1.5 1.548 1.155 1.164 1.166 1.199 1.149 1.105
2 1.517 1.421 1.436 1.435 1.462 1.373 1.354
Avg 1.593 0.986 0.994 0.993 1.016 0.989 0.946
IMP($\%$) 40.615 4.057 4.829 4.733 6.890 4.348
0.5 1.822 0.521 0.526 0.520 0.528 0.558 0.513
1 1.608 0.886 0.895 0.891 0.932 0.917 0.856
180mm 1.5 1.612 1.210 1.222 1.218 1.253 1.203 1.158
2 1.579 1.482 1.497 1.495 1.511 1.438 1.414
Avg 1.655 1.025 1.035 1.031 1.056 1.029 0.985
IMP($\%$) 40.483 3.902 4.831 4.462 6.724 4.276
Figure 5: (a) MAE versus the forecast horizon for all models. (b) CRPS versus the forecast horizon for all models (excluding PER due to poor performance for probabilistic forecasts).
Table 2: Average MAE values (aggregated in 30-min intervals) versus the forecast horizon $h$ for all models, at different locations, namely: E05N, E06, and ASOW6. Bold-faced and underlined values denote the best and second-best performing models, respectively.
Physics-Based Statistical Methods Statistical Deep Learning Methods
Location $h$ (hour) NWP PER ARIMA STGP DeepAR DeepIDE-H DeepMIDE
0.5 1.657 0.477 0.476 0.475 0.478 0.493 0.473
1 1.491 0.811 0.803 0.815 0.831 0.808 0.797
E05N 1.5 1.526 1.103 1.092 1.109 1.123 1.052 1.083
2 1.530 1.360 1.358 1.371 1.383 1.282 1.341
Avg 1.551 0.938 0.932 0.942 0.954 0.909 0.923
IMP($\%$) 40.490 1.599 0.966 2.017 3.250 -1.540
0.5 1.766 0.513 0.524 0.514 0.510 0.534 0.489
1 1.533 0.933 0.956 0.939 0.946 0.880 0.856
E06 1.5 1.498 1.219 1.245 1.234 1.235 1.055 1.089
2 1.459 1.485 1.513 1.506 1.496 1.247 1.315
Avg 1.564 1.037 1.059 1.048 1.047 0.929 0.937
IMP($\%$) 40.090 9.643 11.520 10.592 10.506 -0.861
0.5 1.794 0.542 0.540 0.541 0.565 0.616 0.547
1 1.598 0.847 0.850 0.851 0.927 0.993 0.844
ASOW6 1.5 1.589 1.134 1.150 1.144 1.238 1.328 1.138
2 1.518 1.397 1.413 1.406 1.483 1.572 1.393
Avg 1.625 0.980 0.988 0.986 1.053 1.127 0.980
IMP($\%$) 39.692 0.000 0.810 0.609 6.933 13.044
[…]
@@
Interestingly, this time, DeepIDE-H achieves competitive accuracy compared to DeepMIDE at the central site E06 and downstream site E05N, especially at longer horizons ($h>1$ hour ahead). This suggests that for those locations, the horizontal neighbours provide sufficient spatial context for both models’ horizontal advection kernel, wherein treating height as an independent spatial coordinate (as DeepIDE-H does) is nearly sufficient to capture the local wind field advection. However, this observation does not generalize to other sites. At the upstream boundary location ASOW6, DeepIDE-H performs poorly relative to DeepMIDE. We postulate that horizontal information is scarce, and the forecast depends on propagation between heights, where DeepIDE-H uses a separable vertical kernel, so it cannot redirect information up or down when shear changes. This suggests that the kernel structure in DeepIDE-H, which treats height as an additional spatial coordinate, can capture horizontal advection patterns to some extent, but might not fully capture the complex cross-advection dynamics between different heights.
In terms of all other benchmarks, the location-specific analysis in Table 2 further confirms DeepMIDE’s superiority, showing the lowest average overall MAE across all locations. The maximal improvement appears at E06 (in the range of $\sim$1010-1212% reduction in MAE). We postulate that this may be due to E06 being at a central location, in between the two other spatial sites, making the gain from including neighborhood information greater.
Evaluating probabilistic forecasting performance, Figure 5(b) presents the Continuous Ranked Probability Score (CRPS) for all models across the forecast horizon (We exclude PER due to poor performance for probabilistic forecasts). DeepMIDE consistently achieves the lowest CRPS values, indicating superior probabilistic forecast accuracy compared to all benchmarks. In terms of model adequacy quantified by the $R^{2}$ coefficient, DeepMIDE achieves the highest score ($R^{2}=0.867$), followed by DeepIDE-H ($R^{2}=0.864$), PER ($R^{2}=0.851$), STGP ($R^{2}=0.848$), ARIMA ($R^{2}=0.848$), DeepAR ($R^{2}=0.844$), and NWP ($R^{2}=0.687$). These results further highlight the predictive capabilities of DeepMIDE in both deterministic and probabilistic forecasting.
@@
Figure 6 shows the time series of the one-step wind speed forecasts from DeepMIDE on top of the actual measurements at the hub height of 140140m, during a select time interval in June 2021. The time series of the forecasts show faithful agreement with the actual measurements for all three locations. The corresponding 9595% forecast intervals visually show satisfactory coverage and sharpness. This probabilistic nature of IDE-based models make them well-suited to inform subsequent operational decision-making under uncertainty for offshore wind farms (Pinson2013; papadopoulos2024stochos).
@@
Figure 6: Time series of 11-step ahead forecasts from DeepMIDE on top of the actual wind speed observations at the hub height of 140140m for all three locations (9595% forecast intervals shown).
@@
Wind speed is the major determinant of wind power, and hence, we would like to illustrate the value of our approach for improving wind power prediction accuracy. To do so, we need to convert the wind speed forecasts into the wind power domain. As there are no operational wind farms in the NY/NJ Bight region yet, we leverage an external set of operational data from a wind farm in the United States to construct a statistical wind power curve which will be used for wind-to-power conversion (ding2019data). The power curve is multivariate and it takes two inputs: the hub-height wind speed and the vertical wind shear. The latter reflects the vertical change in wind speed. Let the wind shear be denoted by $S$, then an estimate for wind shear is computed as in (24).
$S=\frac{\log{(Z_{2}/Z_{1})}}{\log{{(b_{1}/b_{2})}}},$ (24)
@@
where $Z_{2}$ and $Z_{1}$ are the wind speeds measured at heights $b_{1}$ and $b_{2}$, respectively (lee2015power). For our data, wind speeds are measured at heights of 100100m, 140140m, and 180180m, with 140140m considered as the hub-height. Consequently, two wind shear values are calculated: an above-hub wind shear ($S_{a}$) using the 180180m/140140m pair, and a below-hub wind shear ($S_{b}$) using the 140140m/100100m pair. Figure 7(a) visualizes the fitted wind power curve (using a multivariate GP model akin to the one proposed by golparvar2021surrogate) with hub height wind speed and above-hub wind shear $S_{a}$. The power output is scaled to the $[0,100]$ interval, where $100$ represents the maximum rated capacity. By applying the constructed power curve, we transform the wind speed observations, along with their forecasts from the six competing models into corresponding wind power values.
@@
Figure 7(b) presents the boxplots of the improvement across all forecasting rolls, horizons, locations, and heights in wind power. DeepMIDE significantly outperforms all its competitors, demonstrating an average improvement of 5.475.47% in wind power forecasting compared to other benchmarks. Enhanced accuracy in wind speed forecasts directly contributes to greater improvements in wind power predictions. Theoretically, wind speed is related to wind power through a cubic relationship, and hence, large improvements in forecasting wind speed are expected to yield even greater improvements in predicting wind power.
@@ 5.2 Discussion
To further elucidate the inner workings of DeepMIDE, we take a closer look at the advection parameter estimates, which are produced using the deep-learning-based functional mapping. Let the norm of the estimated advection vectors at height $g$ and time $t$ denote as $\Lambda_{t}^{g}=\|\boldsymbol{\theta}_{t}^{g}\|$. Figure 8 shows the time series plot of $\Lambda_{t}^{2}$ at height 140140m (corresponding to height index $g=2$ in our study), clearly illustrating a dynamic pattern over time. We zoom in on two consecutive NWP wind speed maps from 21-Jun-2021 02:00:00 to 03:00:00 (orange rectangle, left) and from 04-Jul-2021 05:00:00 to 06:00:00 (green rectangle, right). During the first time interval, the wind speed maps visually show strong advection effects over time, which appears to be captured by the model as reflected by the relatively large advection parameter predictions in the corresponding time series. In contrast, the second time interval shows significantly weaker advection effects (with the two consecutive images being almost static), with correspondingly smaller predicted values in the advection parameter time series. These two examples demonstrate how the deep learning model is able to translate the information from those maps into physically meaningful estimates of advection. Figure 9 shows the boxplots of all $\{\Lambda_{t}^{g}\}_{t=1}^{T}$, partitioned by wind speed (x-axis) and height (different colors), which clearly shows that the values of the predicted advection parameters increase at greater heights and stronger winds—a conclusion which we have observed in our preliminary data analysis of Section 3.3. This further confirms the ability of DeepMIDE to learn key physical insights of the wind field observed from the data.
Figure 8: Time series plot showing the norm of the predicted advection parameters at height 140140m, alongside spatial wind speed images from the numerical model at two consecutive hourly time steps, for two different time intervals: 21-Jun-2021 02:00:00 to 03:00:00 (orange rectangle, left) and from 04-Jul-2021 05:00:00 to 06:00:00 (green rectangle, right).
@@ 6 Conclusions
As the upscaling of offshore wind turbines to unprecedented heights continues, multi-height wind speed forecasts across the now-much-larger rotor area of those ultra-scale generators are needed. In this work, we propose the first attempt for multi-height spatio-temporal wind energy forecasting. Our model, dubbed DeepMIDE, is formulated as a multi-output integro-difference equation model, which fully embraces the full vertical wind profile by incorporating spatial, temporal, and height dependencies, in order to reflect the true complexity of wind interactions coincident with ultra-scale offshore wind turbines. An embedded deep learning architecture directly acts on streams of high-dimensional exogenous weather information in order to encode the physics of local wind field formation in a low-dimensional set of information-rich advection vectors. Those vectors, along with other parameters, are then plugged back into the statistical model for probabilistic space-time forecasting. The proposed approach marries the probabilistic rigor and parsimony of the multi-output integro-difference equation framework, with the representative power of the advanced deep transformer architecture which are used as a “physics extractor” within the statistical model. Evaluated using real-world data from the Northeastern U.S., our model demonstrates significant improvements in both wind speed and power forecasting compared with prevalent forecasting benchmarks.
Future research will investigate more sophisticated modeling structures for the latent process model within the multi-output integro-difference equation framework. This may include handling non-Gaussian data and non-linear dependence, which would obviate the need for data transformations or linear assumptions, but would likely complicate the parameter inference process. Additionally, DeepMIDE could also be extended to model bivariate wind vectors rather than directly acting on wind speed time series. Modeling the wind vector is advantageous because directional transport is central to advection modeling across space and time. It is also beneficial in practice for characterizing wind turbine performance. Methodologically, this requires fundamentally different statistical treatment to jointly model the wind vector components, which are inherently dependent and directionally structured, through a suitable bivariate distribution.
@@
The supplementary materials appended to this manuscript include: (1) Details of the Box-Cox transformation; (2) Details of the architecture and implementation of the deep learning model shown in Figure 4; and (3) Access to codes and data used to reproduce the results of DeepMIDE.
Acknowledgmenets
This work is supported in part by the U.S. National Science Foundation (ECCS-2114422).